Find oxidation of Phosphorous in Phosphoric acid.Can I get a solution to find it’s  oxidation state.

Grade:12th pass

1 Answers

ats
3 years ago
The chemical formula of phosphoric acid is H3PO4.
Let the oxidation state of phosphorous in phosphoric acid be x.
Since it the molecule does not carry any net charge,
                                      +1x3 + x – 2x4 = 0                                [ Oxidation state of H is +1 and oxidation state of O is -2 ]
                                  => 3+x-8 = 0
                                  => x-5 = 0
                                  => x = 5
                                          Hence, the oxidation state of phosphorous in Phosphoric acid is +5.
 
Hope it helps! Pls approve it you found this answer useful.
All the best!